You are given the following cost data:

q TFC TVC
0 25 0
1 25 7
2 25 12
3 25 18
4 25 25
5 25 34
6 25 46
7 25 62
8 25 88

a) If the price of output is $15, how many units of output will this firm produce?

b) What is the total revenue? What is the total cost?

c) Will the firm operate or shut down in the short run? in the long run? Briefly explain your answers.
Answer #1

Answer

TC=TFC+TVC

TC(0)=25+0=25

TC(1)=25+7=32 and so on

MC(n)=TC(n)-TC(n-1)
MC(n)=marginal cost of n th unit
TC(n)=Total cost of n units of output
MC(1)=32-25=7 and so on

TR=P*Q

TR(0)=0*15=0

TR(1)=1*15=15 and so on

Profit=TR-TC

Profit(0)=0-25=-25 and so on

q TFC TVC TC MC TR Profit
0 25 0 25 0 -25
1 25 7 32 7 15 -17
2 25 12 37 5 30 -7
3 25 18 43 6 45 2
4 25 25 50 7 60 10
5 25 34 59 9 75 16
6 25 46 71 12 90 19
7 25 62 87 16 105 18
8 25 88 113 26 120 7

a)

the firm produces at P=MC or the nearest lower MC

MC=12 and P=15 at Q=6 units

so the firm will produce 6 units

============

b)

TR=P*Q=15*6=$150

TR=25+46=71

c)

the firm will operate in the short run as the firm is making a profit and the total revenue is above variable cost

The firm also sustains in the long run if the price is not below the firm's minimum average total cost as the firm is making profit in the short run so the firm is not going to leave the market in the long run.

It will make zero economic profit in the long run.